OpenStudy (rippedblue42):

Where are the asymptotes of f(x)=tan(4x-pi) from x=0 to x=pi/2?

OpenStudy (rippedblue42):

Really need help!!! Will fan and metal!!!

OpenStudy (sshayer):

\[4x-\pi=\frac{ \pi }{ 2 },4x=\pi+\frac{ \pi }{ 2 }=\frac{ 3 \pi }{ 2 },x=\frac{ 3 \pi }{ 8 }\]

OpenStudy (dinamix):